OpenStudy (anonymous):

Hot water is mixed with an amount of ice having an equal mass to that of the water and an initial temperature of 0 °C. What should the initial temperature of the hot water be to achieve a final water temperature of 5 °C with all the ice melted? The specific heat capacity of water is 4.2 kJ/(kg K) and the specific latent heat of fusion for ice is 334 kJ/kg. Assume that no heat is lost to the surroundings.

OpenStudy (anonymous):

hey, that's physics

OpenStudy (anonymous):

so?

OpenStudy (anonymous):

the physics study group is DEAD.

OpenStudy (anonymous):

Although I love physics, I still get lost in it ^^"

OpenStudy (anonymous):

Okay, maybe I can help out with the steps?

OpenStudy (anonymous):

first set the given on one side and RTF (required to find ) on the other side then write down the formulas you know about Heat Capacity

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I'll write them down on paper first.

OpenStudy (anonymous):

wait , there's something wrong in the question , you said that Tinitial = 0 C and you want to find it again?

OpenStudy (anonymous):

no, read the question carefully, the initial temp is for the ice, not water

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

I think I've got the relationship

OpenStudy (anonymous):

Since the mass of ice = the mass of water then M (ice) = M (water) , For ice : Q = mL For water : Q = mC(Tf - Ti ) so solve for m : M (ice) = Q/L M(water) = Q/C(Tf-Ti) since M(ice) = M(water), then: Q/L = Q/C(Tf-Ti) did you end up with this relationship andy?

OpenStudy (anonymous):

something like that, but i think you made it more clear, i get it better now

OpenStudy (anonymous):

glad I could help :) good luck!

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np :)